You plan to work for 40 years and then retire using a 25-year annuity. You want to arrange a retirement income of $4700 per month. You have access to an account that pays an APR of 4.8% compounded monthly.

What size nest egg do you need to achieve the desired monthly yield? (Round your answer to the nearest cent.)
Answer #1

Formula for present value of annuity:

11- (1+i) - Present value = Monthly cash flow x

Interest per month is (0.048/12) = 0.004.

Total number of months is 25*12 = 300.

Present value = 4,700 x 1- (1 + 0.004) -300 0.004

11 - 0.3019160197 Present value = 4,700 x 0.004

Present value = 4, 700 x 174.5209951

Present value = 820, 248.68

The size of nest egg must be $820,248.68.